# Martian time-series unraveled: A multi-scale nested approach with factorial variational autoencoders

### v4

@@ IV Unsupervised source separation
The loss in 9a ensures consistency between the statistics of the recovered signal $\mathbf{s}_{1}$ with the statistics of the observed signals $\mathbf{s}^{i}_{1}$. The loss in 9b ensures that the reconstruction $\mathbf{x}-\mathbf{s}_{1}$ of $\mathbf{n}$ has implicitly the correct statistics by ensuring $\mathbf{n}+\mathbf{s}^{i}_{1}$ has consistent statistics with $\mathbf{x}$. Finally, the loss in 9c promotes statistical independence between the recovered source $\mathbf{s}_{1}$ and the recovered noise $\mathbf{n}$. To facilitate the optimization and to avoid having to choose weighting parameters, each loss term is normalized with respect to the standard deviation of each coefficient in $\Psi_{k}$. In this manner, $\sigma^{2}\left(\Psi_{k}(\mathbf{s}_{1}^{i})\right)$ represents the vector of variances for each coefficient in $\Psi_{k}$ computed across different realizations $\mathbf{s}_{1}^{i}$. The same applies to $\sigma^{2}\left(\Psi_{k}(\mathbf{x}+\mathbf{s}_{1}^{i})\right)$ and $\sigma^{2}\left(\Psi_{k}(\mathbf{s}_{1}^{i},\mathbf{x})\right)$. We then sum the normalized loss terms to define the reconstruction $\mathbf{\widetilde{s}}_{1}$ as the solution to the optimization problem:
$\mathbf{\widetilde{s}}_{1}:=\operatorname*{arg\,min}_{\mathbf{s}_{1}}\,\Big{[}\mathcal{L}_{\text{data}}(\mathbf{s}_{1})+\mathcal{L}_{\text{prior}}(\mathbf{s}_{1})+\mathcal{L}_{\text{cross}}(\mathbf{s}_{1})\Big{]}.$ (10)
The algorithm relies solely on constraints in the scattering spectra space, except for the initialization $\mathbf{\widetilde{s}}_{1}=\mathbf{x}$, which contains valuable information about the mixture of sources in the time domain.
@@ V Results
We demonstrate the effectiveness of our approach on seismic data recorded by the InSight mission. The data consists of seismic records from the SEIS instrument, which has been operational on Mars since 2018. The dataset contains a variety of signals, including marsquakes, glitches, and wind imprints, each with different timescales. We aim to cluster and separate these sources using our proposed framework. We present the results of our approach in the following subsections. Additional results regarding this experiment are provided in appendices A, B and C of the Appendix. Finally, we also showcase our approach in a controlled setting through a stylized example that mimics some of the sources observed in the Mars dataset (cf. appendix D of the Appendix).
@@ V-C Training details of the factorial variational autoencoder
We utilized the entire data recorded during the InSight mission [63] to train the fVAE. During training, we randomly reserved $10\%$ of the data as validation data for tuning a set of hyperparameters. Following the approach of [64], we employed nine clusters at each timescale (see appendix A of the Appendix for a discussion on correspondence with results in [64]). The fVAE was trained using the architecture outlined above, employing a hidden dimension of 1024 and a latent dimension of 32. The Adam optimization algorithm [65] was employed with a learning rate of $10^{-3}$. Training was conducted for $1000$ epochs, utilizing a batch size of $16384$. To address non-differentiability concerns associated with the categorical distribution learning, we utilized the Gumbel-Softmax distribution, enabling a differentiable approximate sampling mechanism for categorical variables [39]. The initial temperature parameter for the Gumbel-Softmax distribution was set to $1.0$, and we exponentially decayed the temperature to a minimum value of $0.5$. Training takes approximately 51 hours on a Tesla V100. We repeated the training with multiple different random seeds and observed relatively good stability. Specifically, the main clusters are still present in different runs, albeit slight changes to the histogram exist (possibly due to the data points lying on the border of different mixture components). We present the reconstruction quality of the pyramidal scattering spectra by the fVAE in appendix B of the Appendix.
@@ V-D Identification of clusters across different timescales
While further investigation will be necessary to fully uncover the nature of the different clusters detected by our study, the results we obtained already demonstrate that the multi-scale approach successfully distinguishes the various phenomena observed in the Martian data. We provide information on all identified clusters within the four timescales considered in appendix A of the Appendix and present the notable clusters in this section. Specifically, Figs. 2 and 3 illustrate two noteworthy identified clusters for the finer ($51.2$ seconds and $3.4$ minutes) and coarser ($13.6$ minutes and $54.6$ minutes) timescales, respectively. To visualize clusters across the four timescales, we depict both the cluster occurrence time histogram along with ten aligned waveforms from that cluster. The cluster occurrence time histogram, i.e., the distribution of windows belonging to that cluster throughout the entire mission, will allow us to identify what times of the day a cluster may exhibit more occurrence, which in turn might reveal clusters that contain atmosphere-surface interaction signals. The aligned waveforms, on the other hand, are useful in identifying the main feature of each cluster. We adopt the approach outlined in [64] to generate the aligned waveforms. This procedure entails calculating, for each cluster within each timescale, the Pearson correlation coefficient between the most probable three component waveform assigned to that cluster (identified via the fVAE) and the remaining three components waveforms in that cluster, adjusted for lag to maximize correlation. Finally, we present the most probable waveform with each component plotted separately, along with nine other aligned waveforms (based on the lag producing the maximum correlation) for the waveforms exhibiting the highest correlation.
(a) Cluster 6 at $51.2$-second timescale
(b) Cluster 7 at $51.2$-second timescale
(c) Cluster 6 at $3.4$-minute timescale
(d) Cluster 8 at $3.4$-minute timescale
@@
(a) Cluster 2 at $13.6$-minute timescale
(b) Cluster 6 at $13.6$-minute timescale
(c) Cluster 0 at $54.6$-minute timescale
(d) Cluster 1 at $54.6$-minute timescale
@@
(a) Cluster 6 at $51.2$-second timescale
(b) Cluster 7 at $51.2$-second timescale
(c) Cluster 6 at $3.4$-minute timescale
(d) Cluster 8 at $3.4$-minute timescale
(e) Cluster 2 at $13.6$-minute timescale
(f) Cluster 6 at $13.6$-minute timescale
(g) Cluster 0 at $54.6$-minute timescale
(h) Cluster 1 at $54.6$-minute timescale
@@ V-E Seasonal impacts on surface-atmosphere interactions
(a) Raw waveform
(b) After separation
(c) Separated signal
(d) Raw waveform close-up
(e) After separation close-up
(f) Separated signal close-up
(g) Raw waveform close-up
(h) After separation close-up
(i) Separated signal close-up
(j) Raw waveform close-up
(k) After separation close-up
(l) Separated signal close-up
Figure 5: Separating glitches from a waveform extracted from cluster $4$ from the $54.6$-minute timescale. The cluster from which the raw waveform is selected is mostly concentrated during the Martian daytime waveform, we select as prior information $300$ samples from cluster $5$ in the $51.2$-second timescale. This cluster has occurrence time histogram mostly concentrated during Martian night and its characteristic waveforms contain no glitches. As such, samples from this cluster would be good candidates for “backgrounds” signals that could allow for separating glitches.
@@ V-F Unsupervised separation of prominent sources
Given access to clusters of glitches of wind-burst noises, in this section, we show that using samples from these clusters, we can perform source separation using the method described in Section IV. In this method, a particular source of interest can be separated from a given time series provided that we have access to several data snippets that does not contain that source of interest. Here explore separating glitches and the imprint of wind bursts. Figs. 5 and 6 summarizes the results, and we provide details and observations for each experiment below. We also provide an example for separating background noise and glitches from a marsquake (cf. appendix C of the Appendix), which extends the application of our method to sources that are less prominent in the dataset.
(a) Raw waveform
(b) After separation
(c) Separated signal
(d) Raw waveform close-up
(e) After separation close-up
(f) Separated signal close-up
(g) Raw waveform close-up
(h) After separation close-up
(i) Separated signal close-up
(j) Raw waveform close-up
(k) After separation close-up
(l) Separated signal close-up
Figure 6: Separating imprints of the wind from a waveform extracted from cluster $3$ from the $54.6$-minute timescale. The cluster from which the raw waveform is selected is mostly concentrated during the Martian daytime and in order to remove glitches from this daytime waveform, we select as prior information $300$ samples from clusters $1$ and $6$ in the $51.2$-second timescale. These clusters have occurrence time histograms mostly concentrated during the day and their characteristic waveform exhibit no wind burst signatures. As such, samples from these clusters would be good candidates for “backgrounds” signals that could allow for separating wind burst related noises.
@@
For the glitch separation experiment, we select a $54.6$-minute long waveform during Martian nighttime with several glitches (cf. Fig. 5(a)). We use $300$ samples from cluster $5$ in the $51.2$-second timescale as its occurrence time histogram mostly concentrated during Martian night and its characteristic waveforms contain no glitches. Figs. 5(b) and 5(c) depicts the waveform after separation of glitches and the separated glitches, respectively. The last four rows of Fig. 5 illustrate the zoomed-in views of the results with columns from left to right being raw waveform, waveform after separation of glitches, and the separated glitch, respectively. These figures indicate the successful separation of glitches as the one-sided pulses are removed with minimal changes to the rest of the waveform. Specifically, the last row of Fig. 5(j) indicates a zoomed-in view of a portion of the raw waveform that contains no glitches, and our method does not separate any coherent signal from the waveform, which is a crucial property of our method.
(a) Broadband events at $51.2$-second timescale
(b) Pressure drops at $51.2$-second timescale
(c) Broadband events at $3.4$-minute timescale
(d) Pressure drops at $3.4$-minute timescale
(e) Broadband events at $13.6$-minute timescale
(f) Pressure drops at $13.6$-minute timescale
(g) Broadband events at $54.6$-minute timescale
(h) Pressure drops at $54.6$-minute timescale
@@
For the wind burst imprint separation we select a Martian daytime $54.6$-minute long waveform (cf. Fig. 6(a)). We use $300$ samples from cluster $1$ in the $51.2$-second timescale as its occurrence time histogram mostly concentrated during Martian day and its characteristic waveforms contain no visible wind bursts. Figs. 6(b) and 6(c) depicts the waveform after separation of glitches and the separated glitches, respectively. The last four rows of Fig. 6 illustrate the zoomed-in views of the results with columns from left to right being raw waveform, waveform after separation of the wind burst imprint, and the separated wave burst imprint, respectively. These figures indicate the successful separation of wind burst imprints as the wind burst imprint (sharp onset followed by ringing oscillations) is removed with minimal changes to the rest of the waveform.

### v5

@@ IV Unsupervised source separation
The loss in 10a ensures consistency between the statistics of the recovered signal $\mathbf{s}_{1}$ with the statistics of the observed signals $\mathbf{s}^{i}_{1}$. The loss in 10b ensures that the reconstruction $\mathbf{x}-\mathbf{s}_{1}$ of $\mathbf{n}$ has implicitly the correct statistics by ensuring $\mathbf{n}+\mathbf{s}^{i}_{1}$ has consistent statistics with $\mathbf{x}$. Finally, the loss in 10c promotes statistical independence between the recovered source $\mathbf{s}_{1}$ and the recovered noise $\mathbf{n}$. To facilitate the optimization and to avoid having to choose weighting parameters, each loss term is normalized with respect to the standard deviation of each coefficient in $\Psi_{k}$. In this manner, $\sigma^{2}\left(\Psi_{k}(\mathbf{s}_{1}^{i})\right)$ represents the vector of variances for each coefficient in $\Psi_{k}$ computed across different realizations $\mathbf{s}_{1}^{i}$. The same applies to $\sigma^{2}\left(\Psi_{k}(\mathbf{x}+\mathbf{s}_{1}^{i})\right)$ and $\sigma^{2}\left(\Psi_{k}(\mathbf{s}_{1}^{i},\mathbf{x})\right)$. We then sum the normalized loss terms to define the reconstruction $\mathbf{\widetilde{s}}_{1}$ as the solution to the optimization problem:
$\mathbf{\widetilde{s}}_{1}=\operatorname*{arg\,min}_{\mathbf{s}_{1}}\,\Big[\mathcal{L}_{\text{data}}(\mathbf{s}_{1})+\mathcal{L}_{\text{prior}}(\mathbf{s}_{1})+\mathcal{L}_{\text{cross}}(\mathbf{s}_{1})\Big].$ (11)
The algorithm relies solely on constraints in the scattering spectra space, except for the initialization $\mathbf{\widetilde{s}}_{1}=\mathbf{x}$, which contains valuable information about the mixture of sources in the time domain. We describe the optimal value that the objective function in 11 converges to in appendix C of the Appendix, which can serve as an approximate quantitative indicator of the quality of the source separation.
@@ V Results
We demonstrate the effectiveness of our approach on seismic data recorded by the InSight mission. The data consists of seismic records from the SEIS instrument, which has been operational on Mars since 2018. The dataset contains a variety of signals, including marsquakes, glitches, and wind imprints, each with different timescales. We aim to cluster and separate these sources using our proposed framework. We present the results of our approach in the following subsections. Additional results regarding this experiment are provided in appendices D, E and H of the Appendix. Finally, we also showcase our approach in a controlled setting through a stylized example that mimics some of the sources observed in the Mars dataset (cf. appendix I of the Appendix).
@@ V-C Training details of the factorial variational autoencoder
We utilized the entire data recorded during the InSight mission [63] to train the fVAE. During training, we randomly reserved $10\%$ of the data as validation data for tuning a set of hyperparameters. Following the approach of [64], we employed nine clusters at each timescale (see appendix D of the Appendix for a discussion on correspondence with results in [64]). Key hyperparameters—such as the number of clusters, dimensionality of the latent space, learning rate, batch size, and network architecture—were selected based on a combination of prior knowledge, grid search, and empirical validation. Specifically, the fVAE was trained using the architecture outlined above, employing a hidden dimension of 1024 and a latent dimension of 32. The Adam optimization algorithm [65] was employed with a learning rate of $10^{-3}$, chosen via grid search for stability. A batch size of $16384$ was used to maximize parallelism within the limits of GPU memory. Network architecture choices were guided by the complexity of the input features and validated through the accuracy in reconstructing the input pyramidal scattering spectra (cf. Figs. 16, 17, 18 and 19 in appendix E of the Appendix), while avoiding overfitting. To address non-differentiability concerns associated with the categorical distribution learning, we utilized the Gumbel-Softmax distribution, enabling a differentiable approximate sampling mechanism for categorical variables [39]. The initial temperature parameter for the Gumbel-Softmax distribution was set to $1.0$, and we exponentially decayed the temperature to a minimum value of $0.5$. Training takes approximately 51 hours on a Tesla V100 GPU, typically converging within this time frame as indicated by the plateauing of the ELBO and stabilization of cluster assignments. We repeated the training with multiple different random seeds and observed relatively good stability; specifically, the main clusters remained consistent across runs, though minor histogram variations were present, possibly due to data points lying near the boundaries of different mixture components.
@@ V-D Identification of clusters across different timescales
While further investigation will be necessary to fully uncover the nature of the different clusters detected by our study, the results we obtained already demonstrate that the multi-scale approach successfully distinguishes the various phenomena observed in the Martian data. We provide information on all identified clusters within the four timescales considered in appendix D of the Appendix and present the notable clusters in this section. Specifically, Figs. 2 and 3 illustrate two noteworthy identified clusters for the finer ($51.2$ seconds and $3.4$ minutes) and coarser ($13.6$ minutes and $54.6$ minutes) timescales, respectively. To visualize clusters across the four timescales, we depict both the cluster occurrence time histogram along with ten aligned waveforms from that cluster. The cluster occurrence time histogram, i.e., the distribution of windows belonging to that cluster throughout the entire mission, will allow us to identify what times of the day a cluster may exhibit more occurrence, which in turn might reveal clusters that contain atmosphere-surface interaction signals. The aligned waveforms, on the other hand, are useful in identifying the main feature of each cluster. We adopt the approach outlined in [64] to generate the aligned waveforms. This procedure entails calculating, for each cluster within each timescale, the Pearson correlation coefficient between the most probable three component waveform assigned to that cluster (identified via the fVAE) and the remaining three components waveforms in that cluster, adjusted for lag to maximize correlation. Finally, we present the most probable waveform with each component plotted separately, along with nine other aligned waveforms (based on the lag producing the maximum correlation) for the waveforms exhibiting the highest correlation.
(a) Glitches without precursor (51.2-second timescale)
(b) Glitches with precursor (51.2-second timescale)
(c) Oscillatory signals (3.4-minute timescale)
(d) High-frequency wind bursts (3.4-minute timescale)
@@
(a) Wind gusts with ringing oscillations (13.6-minute timescale)
(b) Dust devil signatures (13.6-minute timescale)
(c) Sustained wind events (54.6-minute timescale)
(d) Sunrise atmospheric interactions (54.6-minute timescale)
@@
(a) Glitches without precursor (51.2s timescale)
(b) Glitches with precursor (51.2s timescale)
(c) Oscillatory signals (3.4m timescale)
(d) High-frequency wind bursts (3.4m timescale)
(e) Wind gusts with ringing (13.6m timescale)
(f) Dust devil signatures (13.6m timescale)
(g) Sustained wind events (54.6m timescale)
(h) Sunrise-related interactions (54.6m timescale)
@@ V-E Seasonal impacts on surface-atmosphere interactions
(a) Raw mixed signal with glitches
(b) Background signal after glitch separation
(c) Extracted glitch signals
(d) Multiple glitches (raw)
(e) Background after separation
(f) Separated multiple glitches
(g) Glitch with precursor (raw)
(h) Background after separation
(i) Separated glitch with precursor
(j) Clean background region (raw)
(k) Preserved background signal
(l) No spurious separation
Figure 5: Separating glitches from a waveform extracted from cluster 44 from the $54.6$-minute timescale. The cluster from which the raw waveform is selected is mostly concentrated during the Martian daytime waveform, we select as prior information $300$ samples from cluster $5$ in the $51.2$-second timescale. This cluster has occurrence time histogram mostly concentrated during Martian night and its characteristic waveforms contain no glitches. As such, samples from this cluster would be good candidates for “backgrounds” signals that could allow for separating glitches. For comparison with the state-of-practice baseline glitch separation algorithm, see Fig. 20 in appendix F of the Appendix. Refer to Fig. 21 in appendix G of the Appendix for a comparison with the ICA-based source separation method.
@@ V-F Unsupervised separation of prominent sources
Given access to clusters of glitches of wind-burst noises, in this section, we show that using samples from these clusters, we can perform source separation using the method described in Section IV. In this method, a particular source of interest can be separated from a given time series provided that we have access to several data snippets that does not contain that source of interest. Here explore separating glitches and the imprint of wind bursts. Figs. 5 and 6 summarizes the results, and we provide details and observations for each experiment below. We also provide an example for separating background noise and glitches from a marsquake (cf. appendix H of the Appendix), which extends the application of our method to sources that are less prominent in the dataset. Since our approach allows source separation to be performed at a particular time scale, this optimization problem often involves solving multiple independent optimization problems (one for each window of data). To utilize this embarrassingly parallel structure, we run the optimization in parallel on four Tesla V10 devices, for 1000 iterations of the L-BFGS optimization algorithm [68] and a relative termination tolerance of 0.1. The results of glitch separation using the state-of-practice baseline method [69] are presented in appendix F of the Appendix. We also provide a comparison with the ICA-based source separation method in appendix G of the Appendix.
(a) Raw signal with wind bursts
(b) Background after wind separation
(c) Extracted wind burst signals
(d) Wind burst with ringing (raw)
(e) Background after separation
(f) Separated wind burst
(g) Sharp wind onset (raw)
(h) Background after separation
(i) Separated wind signature
(j) Complex wind pattern (raw)
(k) Background after separation
(l) Separated wind pattern
Figure 6: Separating imprints of the wind from a waveform extracted from cluster 33 from the $54.6$-minute timescale. The cluster from which the raw waveform is selected is mostly concentrated during the Martian daytime and in order to remove glitches from this daytime waveform, we select as prior information $300$ samples from clusters $1$ and 66 in the $51.2$-second timescale. These clusters have occurrence time histograms mostly concentrated during the day and their characteristic waveform exhibit no wind burst signatures. As such, samples from these clusters would be good candidates for “backgrounds” signals that could allow for separating wind burst related noises. Refer to Fig. 22 in appendix G of the Appendix for a comparison with the ICA-based source separation method.
@@
For the glitch separation experiment, we select a $54.6$-minute long waveform during Martian nighttime with several glitches (cf. Fig. 5(a)). We use $300$ samples from cluster $5$ in the $51.2$-second timescale as its occurrence time histogram mostly concentrated during Martian night and its characteristic waveforms contain no glitches. The optimization took 104 minutes to complete. Figs. 5(b) and 5(c) depicts the waveform after separation of glitches and the separated glitches, respectively. The last four rows of Fig. 5 illustrate the zoomed-in views of the results with columns from left to right being raw waveform, waveform after separation of glitches, and the separated glitch, respectively. These figures indicate the successful separation of glitches as the one-sided pulses are removed with minimal changes to the rest of the waveform. Specifically, the last row of Fig. 5(j) indicates a zoomed-in view of a portion of the raw waveform that contains no glitches, and our method does not separate any coherent signal from the waveform, which is a crucial property of our method.
(a) Broadband events (quality A) — finest scale
(b) Atmospheric pressure drops — finest scale
(c) Broadband events (quality A) — intermediate scale
(d) Atmospheric pressure drops — intermediate scale
(e) Broadband events (quality A) — coarse scale
(f) Atmospheric pressure drops — coarse scale
(g) Broadband events (quality A) — coarsest scale
(h) Atmospheric pressure drops — coarsest scale
@@
For the wind burst imprint separation we select a Martian daytime $54.6$-minute long waveform (cf. Fig. 6(a)). We use $300$ samples from cluster $1$ in the $51.2$-second timescale as its occurrence time histogram mostly concentrated during Martian day and its characteristic waveforms contain no visible wind bursts. The optimization took 106 minutes to complete. Figs. 6(b) and 6(c) depicts the waveform after separation of glitches and the separated glitches, respectively. The last four rows of Fig. 6 illustrate the zoomed-in views of the results with columns from left to right being raw waveform, waveform after separation of the wind burst imprint, and the separated wave burst imprint, respectively. These figures indicate the successful separation of wind burst imprints as the wind burst imprint (sharp onset followed by ringing oscillations) is removed with minimal changes to the rest of the waveform.
@@ VI Discussion and Conclusion
Our proposed unsupervised, multi-scale clustering and source separation approach demonstrated on Mars InSight data is highly applicable to future planetary missions, including those to Europa and Titan, as these environments present a profound lack of prior knowledge about the types, origins, and timescales of signals that might be encountered. In such unexplored settings, it is impossible to define realistic priors or obtain labeled datasets in advance, which makes unsupervised methods not just advantageous but essential; they can identify and disentangle unknown and potentially novel sources without introducing bias or unrealistic assumptions. This is especially relevant as NASA’s Farside Seismic Suite (FSS) mission will soon deploy a modified version of the SEIS seismometer, originally developed for the InSight mission, to the Moon’s far side where seismic activity remains largely uncharacterized.
While our unsupervised multi-scale clustering and source separation framework is robust to a range of real-world complexities, several fault cases may affect its performance. When multiple sources overlap substantially in both time and frequency, especially if they share similar statistical or spectral characteristics, the clustering and separation steps may yield ambiguous or mixed representations, as the model relies on distinct multi-scale features for differentiation. High levels of additive or non-stationary noise can further obscure source-specific structure, potentially resulting in spurious clusters or reduced separation quality, particularly at finer timescales. Additionally, if the number of clusters per timescale is not well-matched to the true diversity of sources, the model may inadvertently merge distinct sources or fragment single sources across multiple clusters. Such fault cases underscore the importance of careful interpretation and iterative analysis. Importantly, our framework is well-suited for exploratory data analysis: it enables users to iteratively refine clustering and separation, unraveling the underlying sources even in complex or ambiguous scenarios. This iterative process can help identify limitations, guide parameter tuning, and ultimately improve source characterization.